OpenStudy (anonymous):

about trigonometric identities . answer this pls (sinx+cosx)^4 = (1+2sinxcosx)?

OpenStudy (alekos):

have you tried to expand the LHS?

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

pls help me

OpenStudy (alekos):

where did you get this question from? i think there's something missing.

OpenStudy (anonymous):

in my book .

OpenStudy (alekos):

RHS should be (1+2sinxcosx)^2

OpenStudy (anonymous):

oh sorry typos .

OpenStudy (anonymous):

(sinx+cosx)^4 = (1+2sinxcosx)^2

OpenStudy (alekos):

that's what I thought. so if you expand the LHS you get (sinx +cosx)^2 x (sinx +cosx)^2

OpenStudy (alekos):

then we have (sin^2x + 2sinxcosx + cos^2x)(sin^2x + 2sinxcosx + cos^2x) see if you can do the rest?

OpenStudy (anonymous):

oh pls continue

OpenStudy (anonymous):

you used distributive ?

OpenStudy (cp9454):

\[[(sinx +cosx)^2]^2\] \[[\sin^2x+\cos^x + 2sinx cosx]^2\] \[(1+2sinx cosx)^2\]

OpenStudy (anonymous):

then ?

OpenStudy (cp9454):

LHS = RHS , Hence Proved.

OpenStudy (anonymous):

ah what identity did u used ?

OpenStudy (cp9454):

\[\sin^2x + \cos^x = 1\]

OpenStudy (anonymous):

pythagorean . can you pls answer this too 1+sinx/cos x= cos x/1+sinx

OpenStudy (kira_yamato):

As in \[\frac{1 + \sin x}{\cos x} = \frac{\cos x}{1 + \sin x}\]

OpenStudy (anonymous):

yes

OpenStudy (cp9454):

please check your question i think it should be 1-sinx

OpenStudy (anonymous):

okay wait

OpenStudy (cp9454):

\[\frac{(1+\sin x)(1-sinx) }{ cosx(1-sinx) }\] multiplying and dividing by (1-sinx)

OpenStudy (alekos):

Too many typos?

OpenStudy (cp9454):

\[\frac{ 1-\sin^2x }{ cosx(1-sinx) }\] \[\frac{ \cos^2x }{ cosx(1-sinx) }\] \[\frac{ cosx }{ 1-sinx }\]

OpenStudy (alekos):

Well done cp